### v1

@@ 2.2 Size Exclusion Chromatography-Small Angle X-ray Scattering (SEC-SAXS)
Figure 2: Illustration of the effect of DNA bending on the shape of $P(r)$. (A) A group of simple 60 bp DNA models created by Alphafold 3 [ABRAMSOL] three of which were obtained by bending the duplex at base-pair 34. (B) Pair-distance probability distribution functions $P(r)$ for these 4 conformations computed with CRYSOL [SVERGUN995]. ( C/D) the simulated scattering from the 4 duplexes is shown as $\log(I)$ vs. $\log(q)$ (C) or $\log(I)$ vs. $q$ (D). The areas where the computed scattering patterns differ across both types of plots are indicated by the dotted box.
@@ 2.3 Choosing real space instead of reciprocal space for analysis
To perform a comparison of the measured datasets in reciprocal space the data were re-gridded in PRIMUS with a spacing of $\Delta q=0.001\;$Å-1. The re-gridding was carried out using a common $q$-range between all four experiments of 0.00680.49180.0068-0.4918\;Å-1. The ‘data compare’ function in PRIMUS was then used to assess the statistical similarity of the datasets in reciprocal space.
@@ 2.4 Derivation of $P_{\mathrm{exp}}(r)$ from experimental data
The derivation of the pair-distance distribution function $P_{\mathrm{exp}}(r)$ from $I(q)$ versus $q$ via an indirect Fourier transform is performed with the program GNOM of the ATSAS package [MANALASTAS]. The program requires several input parameters which can variably affect the results. Some of these relate to the experimental setup, and others have to be selected. Two of these selected parameters are crucial to the transform process: the maximum diameter of the sample $D_{\mathrm{max}}$ , and the regularisation parameter $\alpha$.
The parameter $D_{\mathrm{max}}$ can be estimated to a good accuracy from the properties of the samples. According to the typical base-pair distance of 3.33.43.3-3.4 Å in the B-form of DNA [SAENGER], a fully rigid ideal 60bp duplex should have a maximum dimension $D_{\mathrm{max}}$ of $\approx$ 205 Å (excluding its hydration shell). Using this as an estimate for what our approximate maximal dimension should be in solution, one of the criteria was that the selection of $D_{\mathrm{max}}$ for each function falls within 10% (184.5 to 225.5 Å) of this value, as there can often be an inherent uncertainty associated with $D_{\mathrm{max}}$ which can be difficult to quantify [TREWHELLA]. As shown by the supplementary figure S4, varying $D_{\mathrm{max}}$ over a range of 2020\;Å (which is very significant in comparison with the length of the DNA sequences) only has a minute effect on the shape of $P_{\mathrm{exp}}(r)$.
@@
For GAGE6 the results of GNOM-5 and GNOM-4 are very close to each other for the shape of $P_{\mathrm{exp}}(r)$ and its estimated error bars. For GAGE6_1 and GAGE6_3 both optimisation procedures converge towards similar shapes although some differences are noticeable. For GAGE6_2, GNOM-5 converges to $\alpha=0.01084$, i.e. a very weak regularisation leading to $P_{\mathrm{exp}}(r)$ showing significant oscillations and large error bars. GNOM-4 converges to $\alpha=1.31$, i.e. a fairly strong regularisation with a very smooth $P_{\mathrm{exp}}(r)$ and small error bars. Figure S5 shows that, for GAGE6_2, the TQE has a wide plateau where it stays almost constant when $\alpha$ varies. This explains why the determination of the optimal $\alpha$ is particularly difficult for this sample. Therefore, for some samples, the calculation of $P_{\mathrm{exp}}(r)$ raises some doubts concerning the validity of the results. However, supplementary figures S5 and S6 show that $\alpha$ essentially determines the magnitude of the humps of $P_{\mathrm{exp}}(r)$ but has little effect on their position with respect to $r$. This is important because Fig. 2 shows that the amplitude of a local DNA bend affects the magnitude of the humps in $P_{\mathrm{exp}}(r)$ but not their position in $r$, which is determined by the location of the bend along the sequence.
@@
Simulated $P(r$) functions (Figure 2) were generated using CRYSOL (Svergun et al., 1995) on atomistic models, applying fake 1% errors to the theoretical intensity and proceeding with the standard derivation of $P(r)$ in Primus.
@@ 2.5.1 General concept
The approach used by Peterlin [PETERLIN] to extract the persistence length of DNA from light scattering data gives some hints on possible approaches for SAXS data. Peterlin introduced a simple polymer model, the Kratky-Porod model, which consists of a chain of rigid segments connected by flexible joints with an energy that depends on the angle $\theta$ between the two joined segments. For this simple model, the theoretical calculation of the scattering function is possible within some limits and its comparison with the experimental data determines the appropriate model parameter, from which the persistence length can be derived. The same idea was refined by Schellmann [SCHELLMAN1974] for a better description of DNA flexibility which took into account dihedral angles $\phi$ between the planes defined by pairs of consecutive segments. However, to allow for analytical calculations, he only considered two limiting cases, either $\phi$ is constrained to zero, i.e. the model evolves in a plane, or $\phi$ is totally random. For DNA the reality lies between these two limits which correspond to either an infinite torsional rigidity ($\phi=0$) or zero torsional rigidity ($\phi$ random, evolving freely without any energy cost). This is why we have chosen an extended Kratky-Porod model with an added energy contribution for the dihedral angles in order to describe not only the persistence length of DNA but also its proper torsional rigidity.
@@ 2.5.2 The DNA polymer model
The DNA polymer model is shown in Fig. 3. It is not concerned with the internal structure of the DNA helix but only with the conformation of its backbone. It consists of $N$ objects representing the base pairs, each separated by the base pair distance in DNA a=3.34a=3.34\;Å. Its N1N-1 segments can be viewed as a generalised Kratky-Porod model. Its conformation is defined by the bending angles $\theta_{n}$, $n=2,\ldots,N-1$ at sites $n$, between segment n1n-1 and $n$ and the dihedral angles $\phi_{n}$, $n=2,\ldots,N-2$, between the planes defined by segments [n1,n][n-1,n] and segments $[n,n+1]$.
The thermal fluctuations of the backbone are controlled by the variation of its bending and torsional energies $E_{\theta}$ and $E_{\phi}$
$H=E_{\theta}+E_{\phi}=\sum_{n=2}^{N-1}K(1-\cos\theta_{n})+\sum_{n=2}^{N-2}C(1-\cos\phi_{n})\;.$ (1)
The constants $K$ and $C$ set the scale of the bending and torsional energies. They are assumed to be the same along the full model, and the bending and twist energies are minimal for $\theta=0$ and $\phi=0$, i.e. the model is homogeneous, with no permanent bending or torsion, in order to avoid any bias towards a specific conformation other that a straight, untwisted, DNA molecule.
Figure 3: The DNA polymer model used in the analysis of the SAXS data.
@@
The constant $C$ determines the torsional rigidity $\xi$ of the model in the linear limit, i.e. for small torsional angles $\phi$ for which $E_{\phi}\approx(1/2)C\phi^{2}$ and a torque $\tau=\partial E_{\phi}/\partial\phi=C\phi$. In this limit a DNA molecule of length $L$ behaves as an elastic rod whose torsional torque grows with torsion as $\tau=(\xi/L)\phi$. For DNA measurements of $\xi$ vary from about $200\;\mathrm{pN\cdot nm^{2}}$ to $480\;\mathrm{pN\cdot nm^{2}}$. Topoisomer distributions of small circles, yielding $\xi\approx 300\;\mathrm{pN\cdot nm^{2}}$ are generally considered as the most reliable measurements [BRYANT]. In the context of our model, for a base pair, i.e. L=a=3.34L=a=3.34\;Å, we get $C=\xi/a=898\,\mathrm{pN\cdot nm}$. When this value, homogeneous to energy (force $\times$ length), is expressed in our energy unit $k_{B}T_{\mathrm{room}}$ we get the dimensionless value $C\approx 218$. Taking into account the broad error bars in the experiments, we have used the value $C=200$ in our calculations as a realistic value to describe the torsional modulus of DNA.
@@
Conformations of this model at room temperature are generated using Monte-Carlo simulations. The first segment $\mathbf{R_{1}}$ is chosen arbitrarily along the $z$ axis of our reference frame, and a second segment $\mathbf{R_{2}}$ is added in the $x,z$ plane by selecting a random value of $\theta_{2}$ in the range $\pi,\pi]$. From two vectors we can generate a local frame $x^{\prime},y^{\prime},z^{\prime}$, using $\mathbf{R_{2}}$ as the $z^{\prime}$ axis, choosing an orthogonal $x^{\prime}$ axis in the $\mathbf{R_{1}}$$\mathbf{R_{2}}$ plane and completing by a third axis $y^{\prime}$ to get an orthonormal frame. A new segment is added in this local frame by selecting $\theta$ in the range $\pi,\pi]$ and the dihedral angle $\phi$ in the range $\pi/2,\pi/2]$, and so on until the $59$ segments corresponding to our $60$-base-pair DNA samples have been added. Each time a new segment has been defined the contribution $E_{\theta}+E_{\phi}$ that it adds to the energy is compared to $k_{B}T$ ($k_{B}T=1$ in our dimensionless units) and accepted or rejected with the Metropolis algorithm [METROPOLIS1953]. Once a conformation of $59$ segments has been completed the probability distribution $P(r)$ of the pair distances between its 60 beads is obtained by a calculation which directly derives from its definition. We compute all pair distances $r_{ij}$, $i=1,\ldots,N$, $j=1,\ldots,N$, $i\not=j$, i.e. N(N1)/2=1770N(N-1)/2=1770 values. Then $P(r)$ is derived from a normalised histogram of these distances by counting how many of them lie within each of $N+1$ boxes of size $a$ centred on the values $r_{k}=(k-\frac{1}{2})a$ (we use real distances, not the dimensionless distances for comparison with experimental data) $P(r_{k})={n_{k}}/{\sum n_{k}}$, where $n_{k}$ is the number of $r_{ij}$ values that fall within box $k$. The least-squares distance $S$ between $P(r)$ and $P_{\mathrm{exp}}(r)$, the normalised distribution function obtained from experimental data, interpolated to get its values at points $r_{k}$, is given by $S=\big{[}\sum_{k}\big{(}P(r_{k})-P_{\mathrm{exp}}(r_{k})\big{)}^{2}\big{]}^{1/2}$.
@@ 2.5.3 Analysis of the saved conformations
We also show the average of $|\theta_{n}(i)|$ restricted to bending angles which exceed a threshold $\theta_{0}=5^{\circ}$ to stress the domains which are prone to large bending.
@@
Figure 4: Statistics of the local bending angles $|\theta_{n}(i)|$ over the 1000 best saved conformations for the GAGE6 sample: the red full line shows $\overline{|\theta_{n}(i)|}$. The black line with circles shows the same average, limited to the bending angles which are above $\theta_{0}=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$
Table 1: Small Angle X-ray scattering data collection parameters
@@
Size Exclusion Chromatograpy SEC-SAS
@@
(B) SAS Data Collection GAGE6 GAGE6_1 GAGE6_2 GAGE6_3
@@
(C) SAS-derived structural parameters GAGE6 GAGE6_1 GAGE6_2 GAGE6_3
@@
From SAS concentration independent method (volume of correlation method) 30.06 31.61 35.53 34.95
@@
SASBDB ID: SASDT72 SASDV87 SASDV97 SASDVA7
However, if the sequence has some intrinsic asymmetry this gives an additional piece of information which can be combined with the SAXS data to solve the dilemma by examining the conformations one by one and reversing some of them so that the peak of large $|\theta_{n}|$ is always on the same side, for instance to the right that we call “forward” conformation. The algorithm to achieve what we call the “orientation” of the conformations is the following. We identify the indices $n_{1}$, $n_{2}$ on both sides of the large $n$ peak, for instance $36$ and $48$ in the case of the GAGE6 sample of Fig. 4. The mirror positions are $n^{\prime}_{1}=N-n_{2}+1$ and $n^{\prime}_{2}=N-n_{1}+1$. Then we scan again the 1000 best conformations and compute for each of them $s=\sum|\theta_{n}|,n=n_{1}\ldots n_{2}$ and $s^{\prime}=\sum|\theta_{n}|,n=n^{\prime}_{1}\ldots n^{\prime}_{2}$. If the difference between $s$ and $s^{\prime}$ is significant, then one of the two peaks is ficticious. If $s>s^{\prime}$ this conformation is already in the forward orientation. If $s^{\prime}>s$ the conformation has to be reversed by switching indices $n$ into $n^{\prime}=N-n+1$. After this check, constructing a plot similar to Fig. 4 with all conformations oriented “forward”, only one of the two peaks remains. As shown in the Results section, this is what happens for the 4 samples that we studied. When reversing a conformation, we also reverse the indices for the dihedral angles $\phi_{n}$ paying attention to the fact that, while $n$ varies from $2$ to N1N-1 for $\theta_{n}$, it varies from $2$ to N2N-2 for $\phi_{n}$ so that, instead of $n^{\prime}$ one has to use $n^{\prime\prime}=N-n$ for $\phi$. Besides the elimination of spurious signals, the orientation process improves the quality of the averaging because it becomes real averaging over the 1000 best configurations instead of the superposition of two sets of data averaged over about 500 conformations.
@@
The analysis can then be extended to detect other properties of the sample. In the Results section, we show how the dihedral angles vary along the sequence, and we also make histograms $H(n,\theta)$ showing, for each site, the distribution of the bending angles among the 1000 best conformations, and similarly for $H(n,\phi)$. These figures are very helpful to get a view of the likely conformations of the samples in solution and to try to understand why some sequences may be favourable for protein binding. Because long polymers, like DNA, fluctuate extensively in solution, a statistical analysis of their most likely conformations is crucial to understanding biological phenomena. It provides information that usefully complements high-resolution structures of protein-DNA complexes, which often only show a snapshot of the final conformational state after binding.

### v2

@@ 2.2 Size Exclusion Chromatography-Small Angle X-ray Scattering (SEC-SAXS)
Figure 2: Illustration of the effect of DNA bending on the shape of $P(r)$. (A) A group of simple 60 bp DNA models, the first of which was created by Alphafold 3 [ABRAMSOL]. Models 2-4 were created by variably bending the first alphafold model manually at base-pair 34. (B) Pair-distance probability distribution functions $P(r)$ for these 4 conformations computed with CRYSOL [SVERGUN995]. ( C/D) the simulated scattering from the 4 duplexes is shown as $\log(I)$ vs. $\log(q)$ (C) or $\log(I)$ vs. $q$ (D). The areas where the computed scattering patterns differ across both types of plots are indicated by the dotted box.
@@ 2.3 Choosing real space instead of reciprocal space for analysis
To perform a comparison of the measured datasets in reciprocal space the data were re-gridded in PRIMUS with a spacing of $\Delta q=0.001\;$Å-1. The re-gridding was carried out using a common $q$-range between all four experiments of 0.00680.49180.0068-0.4918\;Å-1. The ‘data compare’ function in PRIMUS was then used to assess the statistical similarity of the datasets in reciprocal space. To focus on the $q$-range that gives rise to the features that we have observed in $P_{\mathrm{exp}}(r)$, the maximum value of $q$ of the data used for the test was set to 0.180.18\;Å-1. It computes a $p$-value by checking for patterns in the cross-correlation matrix between datasets, which does not depend on the values of the experimental errors [FRANKE2015]. While $p>0.95$ means that two datasets are almost identical, low values of $p$, typically below a threshold of 0.010.050.01-0.05, indicate that the datasets are statistically independent. Such values are generally used to identify outliers, such as sample misloading, denaturing or radiation damage [MANALASTAS]. Intermediate values of $p$ indicate some correlations between the datasets. Supplementary Table 1 shows that the $p$-values obtained by comparing the GAGE6 data successively with data recorded for GAGE6_1, GAGE6_2, GAGE6_3 are 0.493, 0.280, 0.078. This indicates decreasing correlations between the data sets, which are consistent with the nature of our samples. They derive from the same GAGE6 sequence, with only local sequence changes, growing from GAGE6_1 to GAGE6_3. Therefore one should not expect that such samples lead to completely statistically independent datasets as it occurs for outliers due to experimental problems. These $p$ values are interesting because our analysis (see Results section 3) detects a large-bending domain in GAGE6_1 and GAGE6_2 in a portion of the sequence which is not far from the bent domain of detected for GAGE6, which is compatible with a fairly large $p$-value whereas this strongly bent domain disappears for GAGE6_3, which should decrease the correlation with the GAGE6 dataset leading to a lower $p$-value. The statistical analysis appears therefore to provide a very good preview of the results presented in Sec. 3.
@@ 2.4 Derivation of $P_{\mathrm{exp}}(r)$ from experimental data
The derivation of the pair-distance distribution function (Supplementary figure S3) $P_{\mathrm{exp}}(r)$ from $I(q)$ versus $q$ via an indirect Fourier transform is performed with the program GNOM of the ATSAS package [MANALASTAS]. The program requires several input parameters which can variably affect the results. Some of these relate to the experimental setup, and others have to be selected. Two of these selected parameters are crucial to the transform process: the maximum diameter of the sample $D_{\mathrm{max}}$ , and the regularisation parameter $\alpha$.
The parameter $D_{\mathrm{max}}$ can be estimated to a good accuracy from the properties of the samples. According to the typical base-pair distance of 3.33.43.3-3.4 Å in the B-form of DNA [SAENGER], a fully rigid ideal 60bp duplex should have a maximum dimension $D_{\mathrm{max}}$ of $\approx$ 205 Å (excluding its hydration shell). Using this as an estimate for what our approximate maximal dimension should be in solution, one of the criteria was that the selection of $D_{\mathrm{max}}$ for each function falls within 10% (184.5 to 225.5 Å) of this value, as there can often be an inherent uncertainty associated with $D_{\mathrm{max}}$ which can be difficult to quantify [TREWHELLA]. As shown by the supplementary figure S5, varying $D_{\mathrm{max}}$ over a range of 2020\;Å (which is very significant in comparison with the length of the DNA sequences) only has a minute effect on the shape of $P_{\mathrm{exp}}(r)$. Furthermore, as small errors in $D_{\mathrm{max}}$ can alter $P(r)$ slightly [GRANT2015] persistence of features in the respective functions in the face of varying $D_{\mathrm{max}}$ serves as an additional indicator of the robustness of the solution.
@@
For GAGE6 the results of GNOM-5 and GNOM-4 are very close to each other for the shape of $P_{\mathrm{exp}}(r)$ and its estimated error bars. For GAGE6_1 and GAGE6_3 both optimisation procedures converge towards similar shapes although some differences are noticeable. For GAGE6_2, GNOM-5 converges to $\alpha=0.01084$, i.e. a very weak regularisation leading to $P_{\mathrm{exp}}(r)$ showing significant oscillations and large error bars. GNOM-4 converges to $\alpha=1.31$, i.e. a fairly strong regularisation with a very smooth $P_{\mathrm{exp}}(r)$ and small error bars. Figure S6 shows that, for GAGE6_2, the TQE has a wide plateau where it stays almost constant when $\alpha$ varies. This explains why the determination of the optimal $\alpha$ is particularly difficult for this sample. Therefore, for some samples, the calculation of $P_{\mathrm{exp}}(r)$ raises some doubts concerning the validity of the results. However, Supplementary figures S6 and S7 show that $\alpha$ essentially determines the magnitude of the humps of $P_{\mathrm{exp}}(r)$ but has little effect on their position with respect to $r$. This is important because Fig. 2 shows that the amplitude of a local DNA bend affects the magnitude of the humps in $P_{\mathrm{exp}}(r)$ but not their position in $r$, which is determined by the location of the bend along the sequence.
@@
This shows that SAXS data for flexible polymers, which can hardly be distinguished from each other in reciprocal space, even with mathematical tools, can contain detailed information which is revealed when the results are expressed in real space. The above factors together with the simulated $P(r)$ functions which show bending a dsDNA model will cause bumps to appear in the function, the fact that AT tracts have been proven to cause bending of dsDNA, the good fits of each $P(r)$ function to the data (Supplementary Figure S8) and their reasonable TQE (Total Quality Estimate) scores strongly support the validity of the information collected in real space. This is further confirmed by the correlation between the sequence of our DNA samples and the local properties that we detect in our analysis based on $P_{\mathrm{exp}}(r)$.
Simulated $P(r$) functions (Figure 2) were generated using CRYSOL (Svergun et al., 1995) on the first Alphafold 3 model and the 3 subsequent manually bent models, applying fake 1% errors to the theoretical intensity and proceeding with the standard derivation of $P(r)$ in Primus.
@@ 2.5.1 General concept
The approach used by Peterlin [PETERLIN] to extract the persistence length of DNA from light scattering data gives some hints on possible approaches for SAXS data. Peterlin introduced a simple polymer model, the Kratky-Porod model, which consists of a chain of rigid segments connected by flexible joints with an energy that depends on the angle $\theta$ between the two joined segments. For this simple model, the theoretical calculation of the scattering function is possible within some limits and its comparison with the experimental data determines the appropriate model parameter, from which the persistence length can be derived. The same idea was refined by Schellmann [SCHELLMAN1974] for a better description of DNA flexibility which took into account dihedral angles $\varphi$ between the planes defined by pairs of consecutive segments. However, to allow for analytical calculations, he only considered two limiting cases, either $\varphi$ is constrained to zero, i.e. the model evolves in a plane, or $\varphi$ is totally random. For DNA the reality lies between these two limits which correspond to either an infinite torsional rigidity ($\varphi=0$) or zero torsional rigidity ($\varphi$ random, evolving freely without any energy cost). This is why we have chosen an extended Kratky-Porod model with an added energy contribution for the dihedral angles in order to describe not only the persistence length of DNA but also its proper torsional rigidity.
Figure 3: The DNA polymer model used in the analysis of the SAXS data.
@@ 2.5.2 The DNA polymer model
The DNA polymer model is shown in Fig. 3. It is not concerned with the internal structure of the DNA helix but only with the conformation of the line going through the center of the double helix. It consists of $N$ objects representing the base pairs, each separated by the base pair distance in DNA a=3.34a=3.34\;Å. Its N1N-1 segments can be viewed as a generalised Kratky-Porod model. Its conformation is defined by the bending angles $\theta_{n}$, $n=2,\ldots,N-1$ at sites $n$, between segment n1n-1 and $n$ ($\pi<\theta_{n}\leq\pi$) and the dihedral angles $\varphi_{n}$, $n=2,\ldots,N-2$, between the planes defined by segments [n1,n][n-1,n] and segments $[n,n+1]$ ($\pi/2<\varphi_{n}\leq\pi/2$). The reference state for this model is a straight DNA helix, with 11 base pairs per turn, i.e. an angle of about $33^{\circ}$ between them. The $\theta$ and $\varphi$ angles describe the deviations with respect to this structure.
The thermal fluctuations of the backbone are controlled by the variation of its bending and torsional energies $E_{\theta}$ and $E_{\varphi}$
$H=E_{\theta}+E_{\varphi}=\sum_{n=2}^{N-1}K(1-\cos\theta_{n})+\sum_{n=2}^{N-2}C(1-\cos\varphi_{n})\;.$ (1)
The constants $K$ and $C$ set the scale of the bending and torsional energies. They are assumed to be the same along the full model, and the bending and twist energies are minimal for $\theta=0$ and $\varphi=0$, i.e. the model is homogeneous, with no permanent bending or torsion, in order to avoid any bias towards a specific conformation other that a straight, untwisted, DNA molecule.
@@
The constant $C$ determines the torsional rigidity $\xi$ of the model in the linear limit, i.e. for small torsional angles $\varphi$ for which $E_{\varphi}\approx(1/2)C\varphi^{2}$ and a torque $\tau=\partial E_{\varphi}/\partial\varphi=C\varphi$. In this limit a DNA molecule of length $L$ behaves as an elastic rod whose torsional torque grows with torsion as $\tau=(\xi/L)\varphi$. For DNA measurements of $\xi$ vary from about $200\;\mathrm{pN\cdot nm^{2}}$ to $480\;\mathrm{pN\cdot nm^{2}}$. Topoisomer distributions of small circles, yielding $\xi\approx 300\;\mathrm{pN\cdot nm^{2}}$ are generally considered as the most reliable measurements [BRYANT]. In the context of our model, for a base pair, i.e. L=a=3.34L=a=3.34\;Å, we get $C=\xi/a=898\,\mathrm{pN\cdot nm}$. When this value, homogeneous to energy (force $\times$ length), is expressed in our energy unit $k_{B}T_{\mathrm{room}}$ we get the dimensionless value $C\approx 218$. Taking into account the broad error bars in the experiments, we have used the value $C=200$ in our calculations as a realistic value to describe the torsional modulus of DNA.
@@
Conformations of this model at room temperature are generated using Monte-Carlo simulations. The first segment $\mathbf{R_{1}}$ is chosen arbitrarily along the $z$ axis of our reference frame, and a second segment $\mathbf{R_{2}}$ is added in the $x,z$ plane by selecting a random value of $\theta_{2}$ in the range $\pi<\theta_{2}\leq\pi$. From two vectors we can generate a local frame $x^{\prime},y^{\prime},z^{\prime}$, using $\mathbf{R_{2}}$ as the $z^{\prime}$ axis, choosing an orthogonal $x^{\prime}$ axis in the $\mathbf{R_{1}}$$\mathbf{R_{2}}$ plane and completing by a third axis $y^{\prime}$ to get an orthonormal frame. A new segment is added in this local frame by selecting $\theta$ in the range $\pi<\theta\leq\pi$ and the dihedral angle $\varphi$ in the range $\pi/2<\varphi\leq\pi/2$, and so on until the $59$ segments corresponding to our $60$-base-pair DNA samples have been added. Each time a new segment has been defined the contribution $E_{\theta}+E_{\varphi}$ that it adds to the energy is compared to $k_{B}T$ ($k_{B}T=1$ in our dimensionless units) and accepted or rejected with the Metropolis algorithm [METROPOLIS1953]. Once a conformation of $59$ segments has been completed the probability distribution $P(r)$ of the pair distances between its 60 beads is obtained by a calculation which directly derives from its definition. We compute all pair distances $r_{ij}$, $i=1,\ldots,N$, $j=1,\ldots,N$, $i\not=j$, i.e. N(N1)/2=1770N(N-1)/2=1770 values. Then $P(r)$ is derived from a normalised histogram of these distances by counting how many of them lie within each of $N+1$ boxes of size $a$ centred on the values $r_{k}=(k-\frac{1}{2})a$ (we use real distances, not the dimensionless distances for comparison with experimental data) $P(r_{k})={n_{k}}/{\sum n_{k}}$, where $n_{k}$ is the number of $r_{ij}$ values that fall within box $k$. The least-squares distance $S$ between $P(r)$ and $P_{\mathrm{exp}}(r)$, the normalised distribution function obtained from experimental data, interpolated to get its values at points $r_{k}$, is given by $S=\big[\sum_{k}\big(P(r_{k})-P_{\mathrm{exp}}(r_{k})\big)^{2}\big]^{1/2}$.
@@ 2.5.3 Analysis of the saved conformations
We also show the average of $|\theta_{n}(i)|$ restricted to bending angles which exceed a threshold $\theta_{0}=5^{\circ}$ to stress the domains which are prone to large bending. The average diehdral angle $\overline{|\varphi_{n}(i)|}$ and the standard deviation $\sigma_{\varphi}$ are computed with equations similar to Eqs. (2) and (3).
@@
Table 1: Sample information and characterization - Small Angle X-Ray Scattering parameters
@@
Size Exclusion Chromatograpy SEC-SAXS
@@
(B) SAXS Data Collection GAGE6 GAGE6_1 GAGE6_2 GAGE6_3
@@
(C) SAXS-derived structural parameters GAGE6 GAGE6_1 GAGE6_2 GAGE6_3
@@
From SAXS concentration independent method (volume of correlation method) 30.06 31.61 35.53 34.95
@@
SASBDB ID: SASDXH7 SASDV87 SASDV97 SASDVA7
However, if the sequence has some intrinsic asymmetry this gives an additional piece of information which can be combined with the SAXS data to solve the dilemma by examining the conformations one by one and reversing some of them so that the peak of large $|\theta_{n}|$ is always on the same side, for instance to the right that we call “forward” conformation. The algorithm to achieve what we call the “orientation” of the conformations is the following. We identify the indices $n_{1}$, $n_{2}$ on both sides of the large $n$ peak, for instance $36$ and $48$ in the case of the GAGE6 sample of Fig. 4. The mirror positions are $n^{\prime}_{1}=N-n_{2}+1$ and $n^{\prime}_{2}=N-n_{1}+1$ i.e. $n^{\prime}_{1}=13$ and $n^{\prime}_{2}=25$ for the example above. Then we scan again the 1000 best conformations and compute for each of them $s=\sum|\theta_{n}|,n=n_{1}\ldots n_{2}$ and $s^{\prime}=\sum|\theta_{n}|,n=n^{\prime}_{1}\ldots n^{\prime}_{2}$. If the difference between $s$ and $s^{\prime}$ is significant, which is evaluated by calculating $\Delta s/s=|s^{\prime}-s|/s$, then one of the two peaks is ficticious. If $s>s^{\prime}$ this conformation is already in the forward orientation. If $s^{\prime}>s$ the conformation has to be reversed by switching indices $n$ into $n^{\prime}=N-n+1$. After this check, constructing a plot similar to Fig. 4 with all conformations oriented “forward”, only one of the two peaks remains. As shown in the Results section, this is what happens for the 4 samples that we studied. When reversing a conformation, we also reverse the indices for the dihedral angles $\varphi_{n}$ paying attention to the fact that, while $n$ varies from $2$ to N1N-1 for $\theta_{n}$, it varies from $2$ to N2N-2 for $\varphi_{n}$ so that, instead of $n^{\prime}$ one has to use $n^{\prime\prime}=N-n$ for $\varphi$. Besides the elimination of spurious signals, the orientation process improves the quality of the averaging because it becomes real averaging over the 1000 best configurations instead of the superposition of two sets of data averaged over about 500 conformations.
Figure 4: Statistics of the local bending angles $|\theta_{n}(i)|$ over the 1000 best saved conformations for the GAGE6 sample: the red full line shows $\overline{|\theta_{n}(i)|}$. The black line with circles shows the same average, limited to the bending angles which are above $\theta_{0}=5^{\circ}$, and the dashed blue curve shows the standard deviation of $|\theta_{n}(i)|$. The full brown vertical lines show the positions $n_{1}=36$, $n_{2}=48$ which have been used for the ’orientation’ of the conformations and the dashed brown vertical lines show the mirror positions $n^{\prime}_{1}=13$ and $n^{\prime}_{2}=25$.
@@
The analysis can then be extended to detect other properties of the sample. In the Results section, we show how the dihedral angles vary along the sequence, and we also calculate the probability $P(n,|\theta|)$ showing, for each site, the distribution of the bending angles among the 1000 best conformations, and similarly for $P(n,|\varphi|)$. These figures are very helpful to get a view of the likely conformations of the samples in solution and to try to understand why some sequences may be favourable for protein binding. Because long polymers, like DNA, fluctuate extensively in solution, a statistical analysis of their most likely conformations is crucial to understanding biological phenomena. It provides information that usefully complements high-resolution structures of protein-DNA complexes, which often only show a snapshot of the final conformational state after binding.